### 01 Dec 2023 v1

# and the Celestial Sphere

Abstract
It is shown that in the presence of a nonvanishing cosmological constant, Strominger’s infinite-dimensional $\mathrm{w_{1+\infty}}$ algebra of soft graviton symmetries is modified in a simple way. The deformed algebra contains a subalgebra generating $SO(1,4)$ or $SO(2,3)$ symmetry groups of $\text{dS}_{4}$ or $\text{AdS}_{4}$, depending on the sign of the cosmological constant. The transformation properties of soft gauge symmetry currents under the deformed $\mathrm{w_{1+\infty}}$ are also discussed.
Introduction. — The conservation laws reflect the symmetries of nature and provide a key to understanding the physical universe. What was less appreciated until few years ago is the importance of a rather specialized area of quantum field theory and gravity devoted to studying the physical processes involving “soft” particles, with very low energies. The zero energy limit of the scattering amplitudes involving soft gauge bosons, gravitons and other particles are described by “soft theorems” Weinberg:1965nx . The long wavelengths of soft particles allow probing the large scale structure of universe, particularly the past and future asymptotic infinities. Hence, as shown by Strominger and collaborators Strominger:2013lka ; Strominger:2013jfa ; He:2014laa ; Kapec:2014opa ; Strominger:2014pwa , soft theorems are closely related to the conservation laws and symmetries. Actually, every single (known) soft theorem in asymptotically flat spacetime has been associated with an infinite number of symmetries of the celestial sphere at null infinity. These include Poincaré and extended BMS symmetries. Bondi:1962px ; Sachs:1962wk ; Barnich:2009se . The connection between soft theorems and asymptotic symmetries has laid the foundations for the celestial holography program which aims at describing four-dimensional physics in terms of a two-dimensional conformal field theory on the celestial sphere (CCFT) Raclariu:2021zjz .
Two tears ago, Strominger performed a systematic study of the symmetries associated with soft particles carrying positive helicities Strominger:2021lvk . He showed that all these symmetries are encompassed in an infinite-dimensional $\mathrm{w_{1+\infty}}$ algebra. $\mathrm{w_{1+\infty}}$ was extracted from the algebra of soft currents encoded in the operator product expansion (OPE) of celestial primary operators associated with gravitons and gauge bosons Guevara:2021abz . In CCFT, OPEs can be obtained from the collinear limits of (celestial) scattering amplitudes Fan:2019emx ; Pate:2019lpp . Strominger’s results follow from tree-level amplitudes evaluated in flat spacetime with vanishing cosmological constant. [^1] Various aspects of soft symmetry algebras are discussed in Refs.Ball:2021tmb ; Himwich:2021dau ; Mago:2021wje ; Costello:2022wso ; Fan:2022vbz ; Casali:2022fro ; Bu:2022iak ; Stieberger:2022zyk ; Bittleston:2022jeq ; Monteiro:2022xwq ; Banerjee:2023rni ; Taylor:2023bzj ; Garner:2023izn ; Bittleston:2023bzp ; Melton:2022fsf ; Drozdov:2023qoy ; Lipstein:2023pih ; Costello:2023hmi ; Adamo:2023zeh ; Ball:2023ukj ; Krishna:2023ukw ; Banerjee:2023jne ; Banerjee:2023bni
[…]
The idea. — The idea originates from the recent work of Alday, Hansen and Silva, who computed the amplitude for the scattering of four gravitons on $\mathrm{AdS_{5}\times S_{5}}$ Alday:2022uxp ; Alday:2022xwz ; Alday:2023jdk ; Alday:2023mvu . Although no rigorous definition of the S-matrix exists for non-asymptotically flat spacetimes like AdS, they formally used AdS/CFT correspondence Maldacena:1997re and expanded the amplitude in the inverse curvature radius $R^{-2}\propto\Lambda$. In the limit $\Lambda\to 0$, they obtained the well known Virasoro-Shapiro amplitude. The subleading term is of order ${\cal O}(\Lambda)$. Instead of considering it as a part of a full-fledged AdS S-matrix element, we can consider it as a curvature-induced correction the scattering amplitude in flat spacetime. For comparison, the proton-proton cross sections measured at the LHC also receive similar corrections although protons do not fly in from the cosmological horizon but come from a bottle of hydrogen stored in Meyrin, Switzerland. The subleading term has a very interesting property which becomes transparent after taking the string zero slope limit, while keeping fixed the gravitational coupling constant $\kappa$. Then, for the gravitons in the (++)(-\,-+\,+)( - - + + ), i.e. in the MHV helicity configuration,
$A^{(1)}(s,t,u)=\kappa\Lambda\frac{\langle 12\rangle^{6}[43]}{\langle 13\rangle\langle 14\rangle\langle 23\rangle\langle 24\rangle\langle 34\rangle}\Big{(}\frac{1}{s}+\frac{1}{t}+\frac{1}{u}\Big{)},$ (1)
where $s,t,u$ are the Mandelstam variables and for the momentum spinors, we used the notation of Ref.Taylor:2017sph . We are interested in the limit of collinear (light-like) momenta $p_{3}$ and $p_{4}$, when $s\to 0$. For our purposes, it is convenient to parametrize momenta in terms of light-cone energies $\omega$ and complex coordinates $z$ on the celestial sphere Raclariu:2021zjz . Then, for two arbitrary light-like momenta $p_{i}$ and $p_{j}$,
$\langle ij\rangle=\sqrt{\omega_{i}\omega_{j}}(z_{i}-z_{j})\ ,\quad[ij]=\sqrt{\omega_{i}\omega_{j}}(\bar{z}_{j}-\bar{z}_{i})\ ,\quad 2p_{i}p_{j}=\omega_{i}\omega_{j}|z_{i}-z_{j}|^{2}\ .$ (2)
In order to define the collinear limit, we complexify the momenta, so that $z$ and $\bar{z}$ can be considered as independent complex variables Taylor:2017sph . The combined momentum of the collinear pair is defined as $P=p_{3}+p_{4}$. In the collinear limit $z_{3}\to z_{4}$ (while keeping $\bar{z}_{3}$ and $\bar{z}_{4}$ fixed) and $P^{2}\to 0$. The leading collinear singularity has the form
$A^{(1)}(s,t,u)\approx\kappa\Lambda\frac{\omega_{P}^{2}}{\omega_{3}^{2}\omega_{4}^{2}(z_{3}-z_{4})^{2}}\left(\frac{\langle 12\rangle^{6}}{\langle 1P\rangle^{2}\langle 2P\rangle^{2}}\right),$ (3)
therefore we obtain the three-graviton amplitude (enclosed in the brackets) times the collinear factor with a double pole $(z_{3}-z_{4})^{-2}$. This is a stronger collinear singularity than the single pole encountered in flat spacetime.
Graviton OPEs. — The leading term in the OPE of the primary CCFT operators $G^{+}_{\Delta}(z)$ with dimensions $\Delta$, associated with the (positive helicity) gluons can be extracted from the amplitude (3) in the same way as in Ref.Fan:2019emx . We obtain
$G^{+}_{\Delta_{3}}(z_{3},\bar{z}_{3})\,G^{+}_{\Delta_{4}}(z_{4},\bar{z}_{4})\sim\kappa\Lambda\frac{B(\Delta_{3}-2,\Delta_{4}-2)}{z_{34}^{2}}G^{+}_{\Delta_{3}+\Delta_{4}-2}(z_{4},\bar{z}_{4})\ ,$ (4)
where $z_{34}=z_{3}-z_{4}$. One can also extract a single pole term $\sim z_{34}^{-1}$, which is necessary for the symmetry of the operator product under $3\leftrightarrow 4$. There is, however, a problem with this OPE. If one proceeds along the lines of Ref.Guevara:2021abz , and extracts the algebra of soft currents associated with the graviton operators, it fails Jacobi identity. It is not difficult though, to find a slight modification of the OPE coefficients leading to consistent double and single pole singularities:
$G^{+}_{\Delta_{3}}(z_{3},\bar{z}_{3})G^{+}_{\Delta_{4}}(z_{4},\bar{z}_{4})=$ $-\frac{\kappa}{2}\frac{\bar{z}_{34}}{z_{34}}B(\Delta_{3}-1,\Delta_{4}-1)G^{+}_{\Delta_{3}+\Delta_{4}}(z_{4},\bar{z}_{4})$
$+\frac{\kappa\,\Lambda}{2}\frac{\Delta_{3}+\Delta_{4}}{z_{34}^{2}}B(\Delta_{3}-2,\Delta_{4}-2)G^{+}_{\Delta_{3}+\Delta_{4}-2}(z_{4},\bar{z}_{4})$
$+\frac{\kappa\,\Lambda}{2}\frac{\Delta_{3}}{z_{34}}B(\Delta_{3}-2,\Delta_{4}-2)\partial G^{+}_{\Delta_{3}+\Delta_{4}-2}(z_{4},\bar{z}_{4})\,,$ (5)
where, for completeness, we also included, in the first term, the contribution of the zero slope limit of the Virasoro-Shapiro amplitude in flat spacetime. The coefficient of the double pole term, see the second term on the r.h.s. of Eq.($\mathbf{w_{1+\infty}}$ Algebra with a Cosmological Constant                                    and the Celestial Sphere), contains an extra factor of $(\Delta_{3}+\Delta_{4})/2$ as compared to the collinear limit (4). We can only speculate that it is due to a modified form of the momentum conservation law in curved spacetime. Indeed, as shown below, it will change the commutation relations of the “momentum” operators in a way expected for a spacetime with constant curvature.
From OPEs to cosmological $\mathrm{w_{1+\infty}}$. — After including the antiholomorphic descendants in the OPE of Eq.($\mathbf{w_{1+\infty}}$ Algebra with a Cosmological Constant                                    and the Celestial Sphere), it acquires the form:
$G^{+}_{\Delta_{3}}(z_{3},\bar{z}_{3})G^{+}_{\Delta_{4}}(z_{4},\bar{z}_{4})=$ $-\frac{\kappa}{2}\frac{1}{z_{34}}\sum_{n=0}^{\infty}B(\Delta_{3}-1+n,\Delta_{4}-1)\frac{(\bar{z}_{34})^{n+1}}{n!}\bar{\partial}^{n}G^{+}_{\Delta_{3}+\Delta_{4}}(z_{4},\bar{z}_{4})$
$+\frac{\kappa\,\Lambda}{2}\frac{\Delta_{3}+\Delta_{4}}{z_{34}^{2}}\sum_{n=0}^{\infty}B(\Delta_{3}-2+n,\Delta_{4}-2)\frac{(\bar{z}_{34})^{n}}{n!}\bar{\partial}^{n}G^{+}_{\Delta_{3}+\Delta_{4}-2}(z_{4},\bar{z}_{4})$
$+\frac{\kappa\,\Lambda}{2}\frac{\Delta_{3}}{z_{34}}\sum_{n=0}^{\infty}B(\Delta_{3}-2+n,\Delta_{4}-2)\frac{(\bar{z}_{34})^{n}}{n!}\partial\,\bar{\partial}^{n}G^{+}_{\Delta_{3}+\Delta_{4}-2}(z_{4},\bar{z}_{4})\,.$ (6)
Next, we define the conformally soft graviton operators,
$H^{k}=\lim_{\epsilon\to 0}\epsilon G^{+}_{k+\epsilon}\ ,~{}~{}\qquad k=2,1,0,-1,\dots,$ (7)
with conformal weights $\{h,\bar{h}\}=\{(k+2)/2,(k-2)/2\}$. We represent them as truncated antiholomorphic series,
$H^{k}(z,\bar{z})=\sum_{n=\frac{k-2}{2}}^{\frac{2-k}{2}}\frac{H^{k}_{n}(z)}{\bar{z}^{n+\frac{k-2}{2}}}\ ,$ (8)
and further expand the holomorphic coefficients:
$H^{k}_{n}(z)=\sum_{a=-\infty}^{\infty}\frac{H^{k}_{a,n}}{z^{a+\frac{k+2}{2}}}$ (9)
The OPE of Eq.(6) translates into the following algebra of soft currents:
$[H^{k}_{a,m},H^{l}_{b,n}]=$ $-\frac{\kappa}{2}[n(2-k)-m(2-l)]\frac{\left(\frac{2-k}{2}-m+\frac{2-l}{2}-n-1\right)!\left(\frac{2-k}{2}+m+\frac{2-l}{2}+n-1\right)!}{\left(\frac{2-k}{2}-m\right)!\left(\frac{2-l}{2}-n\right)!\left(\frac{2-k}{2}+m\right)!\left(\frac{2-l}{2}+n\right)!}H^{k+l}_{a+b,m+n}$
$+\frac{\kappa\,\Lambda}{2}\left(l\,a-k\,b\right)\frac{\left(\frac{2-k}{2}-m+\frac{2-l}{2}-n\right)!\left(\frac{2-k}{2}+m+\frac{2-l}{2}+n\right)!}{\left(\frac{2-k}{2}-m\right)!\left(\frac{2-l}{2}-n\right)!\left(\frac{2-k}{2}+m\right)!\left(\frac{2-l}{2}+n\right)!}H^{k+l-2}_{a+b,m+n}\,.$ (10)
To make connection with Strominger’s $\mathrm{w_{1+\infty}}$, we define
$w^{p}_{a,m}=\frac{1}{\kappa}(p-m-1)!(p+m-1)!H^{-2p+4}_{a,m}\,,$ (11)
where $p$ run over the positive half integers
[…]
and the condition of the truncated antiholomorphic mode expansion (8) turns into the following constraint on the indices labelled by $m$:
$1-p\leq m\leq p-1\,.$ (13)
The indices $a$, associated with the holomorphic modes, are integer for integer $p$ and half-integer for half-integer $p$, similar to $m$, but their range is not restricted:
$a=-\infty,\dots,p-1,p,p+1,\dots,\infty.$ (14)
The algebra (10), written in terms of $w$ generators, becomes
$\,\,[w^{p}_{a,m},\,w^{q}_{b,n}]=[m(q-1)-n(p-1)]w^{p+q-2}_{a+b,m+n}-\Lambda[a(q-2)-b(p-2)]w^{p+q-1}_{a+b,m+n}\,.$ (15)
It is easy to check that it satisfies Jacobi identity and closes within the range of indices given in Eqs.(12-14).
Properties of deformed algebra. — In order to understand the structure of the deformed $\mathrm{w_{1+\infty}}$ algebra (15) and the role of the cosmological constant, we note that it contains a closed subalgebra of 10 generators: $w^{1}_{a,0}$ with $a=-1,0,1$, $w^{\frac{3}{2}}_{\pm\frac{1}{2},\pm\frac{1}{2}}$, and $w^{2}_{0,m}$ with $m=-1,0,1$. It reads:
$[w^{1}_{a,0},w^{2}_{0,m}]=0\ ,$ (16)
$[w^{1}_{a,0},w^{1}_{b,0}]=\Lambda(a-b)w^{1}_{a+b,0}\ ,$ (17)
$[w^{2}_{0,m},w^{2}_{0,n}]=(m-n)w^{2}_{0,m+n}\ ,$ (18)
$[w^{1}_{a,0},w^{\frac{3}{2}}_{k,l}]=\Lambda\left(\frac{a}{2}-k\right)w^{\frac{3}{2}}_{a+k,l}\ ,$ (19)
$[w^{2}_{0,m},w^{\frac{3}{2}}_{k,l}]=\left(\frac{m}{2}-l\right)w^{\frac{3}{2}}_{k,m+l}\ ,$ (20)
$[w^{\frac{3}{2}}_{i,j},w^{\frac{3}{2}}_{k,l}]=\frac{1}{2}(j-l)w^{1}_{i+k,j+l}+\frac{\Lambda}{2}(i-k)w^{2}_{i+k,j+l}\ .~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (21)
In Strominger’s $\mathrm{w_{1+\infty}}$, $w^{1}_{a,0}$ are c-number operators. In the present algebra, however, they do not commute. To see it in a more transparent way, we define
$w^{1}_{a,0}=\Lambda\,L_{a}\,,\quad w^{2}_{0,m}=\bar{L}_{m}\,,\quad w^{\frac{3}{2}}_{k,l}=P_{k,l}\,.$ (22)
In terms of these operators, the above algebra reads
$[L_{a},\bar{L}_{m}]=0\ ,$ (23)
$[L_{a},L_{b}]=(a-b)L_{a+b}\,,$ (24)
$[\bar{L}_{m},\bar{L}_{n}]=(m-n)\bar{L}_{m+n}\,,$ (25)
$[L_{a},P_{k,l}]=\left(\frac{a}{2}-k\right)P_{a+k,l}\,,$ (26)
$[\bar{L}_{m},P_{k,l}]=\left(\frac{m}{2}-l\right)P_{k,m+l}\,,$ (27)
$[P_{i,j},P_{k,l}]=\Lambda\,j\,\delta_{j,-l}\,L_{i+k}+\Lambda\,i\,\delta_{i,-k}\,\bar{L}_{j+l}\ .~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}$ (28)
In the limit of $\Lambda=0$, it is the Poincaré subalgebra of extended BMS symmetry, with four translations $P_{\mu}\,,\mu=0,1,2,3$, defined by,
$P_{-\frac{1}{2},-\frac{1}{2}}$ $=P_{0}+P_{3}\,,\quad P_{-\frac{1}{2},\frac{1}{2}}=P_{1}-iP_{2}\,,$
$P_{\frac{1}{2},-\frac{1}{2}}$ $=P_{1}+iP_{2}\,,\quad P_{\frac{1}{2},\frac{1}{2}}=P_{0}-P_{3}\,,$ (29)
and with six Virasoro operators $L_{1,0,1}\,,\bar{L}_{-1,0,1}$ related to the Lorentz generators $M_{\mu\nu}=-M_{\nu\mu}$ in the following way Stieberger:2018onx ; Fotopoulos:2019vac :
$M_{23}+iM_{10}$ $=-L_{-1}+L_{1}\,,\qquad-M_{23}+iM_{10}=-\bar{L}_{-1}+\bar{L}_{1}\,,$ (30)
$M_{20}+iM_{13}$ $=-L_{-1}-L_{1}\,,\qquad-M_{20}+iM_{13}=-\bar{L}_{-1}-\bar{L}_{1}\,,$ (31)
$M_{21}+iM_{03}$ $=-2L_{0}\,,~{}~{}~{}\qquad\quad-M_{21}+iM_{03}=-2\bar{L}_{0}\,,$ (32)
[…]
$P_{\mu}=\frac{\sqrt{|\Lambda|}}{2}M_{\mu 4}\,.$ (33)
Then the algebra of Eqs.(23-28) can be written as
$[M_{\mu\nu},M_{\rho\lambda}]=i(\eta_{\mu\rho}\,M_{\nu\lambda}+\eta_{\nu\lambda}M_{\mu\rho}-\eta_{\nu\rho}M_{\mu\lambda}-\eta_{\mu\lambda}M_{\nu\rho})\,,$ (34)
with
[…]
Gauge theory. — We started this discussion from the collinear limit of curvature corrections to the graviton scattering amplitudes Alday:2022uxp ; Alday:2022xwz ; Alday:2023jdk ; Alday:2023mvu . It would be very interesting to study gauge theories coupled to gravity in a similar way, in particular curvature corrections to graviton-gauge boson interactions in Einstein-Yang-Mills theory. Here, we also expect a double pole in the OPE of the positive helicity graviton operator $G^{+}_{\Delta}(z,\bar{z})$ and the gauge boson operator $O^{+d}_{\Delta}(z,\bar{z})$, where $d$ labels the group index. It would modify the corresponding OPE in the following way:
$G^{+}_{\Delta_{1}}(z_{1},\bar{z}_{1})O^{+d}_{\Delta_{2}}(z_{2},\bar{z}_{2})=$ $-\frac{\kappa}{2}\frac{\bar{z}_{12}}{z_{12}}B(\Delta_{1}-1,\Delta_{2})O^{+d}_{\Delta_{1}+\Delta_{2}}(z_{2},\bar{z}_{2})$
$+\frac{\kappa\Lambda}{2}\,\frac{\Delta_{1}+\Delta_{2}-1}{z_{12}^{2}}\,B(\Delta_{1}-2,\Delta_{2}-1)O^{+d}_{\Delta_{1}+\Delta_{2}-2}(z_{2},\bar{z}_{2})$
$+\frac{\kappa\Lambda}{2}\,\frac{\Delta_{1}}{z_{12}}\,B(\Delta_{1}-2,\Delta_{2}-1)\partial O^{+d}_{\Delta_{1}+\Delta_{2}-2}(z_{2},\bar{z}_{2})\,.$ (36)
To see what is the corresponding deformation of the symmetry algebra, we define the conformally soft gluon operators,
$R^{k,d}=\lim_{\epsilon\rightarrow 0}\epsilon\,O^{+d}_{k+\epsilon},\quad k=1,0,-1,\dots\,$ (37)
with conformal weights $\{h,\bar{h}\}=\{(k+1)/2,(k-1)/2\}$. We represent them as a truncated holomorphic series:
$R^{k,d}(z,\bar{z})=\sum_{n=\frac{k-1}{2}}^{\frac{1-k}{2}}\frac{R^{k,d}_{n}(z)}{\bar{z}^{n+\frac{k-1}{2}}}\ ,$ (38)
and further expand the holomorphic coefficients,
$R^{k,d}_{n}(z)=\sum_{a=-\infty}^{\infty}\frac{R^{k,d}_{a,n}}{z^{a+\frac{k+1}{2}}}\,.$ (39)
The algebra of the soft currents following from the OPE (36) has the form:
$[H^{k}_{a,m},R^{l,d}_{b,n}]=$ $-\frac{\kappa}{2}[n(2-k)-m(1-l)]\frac{\left(\frac{2-k}{2}-m+\frac{1-l}{2}-n-1\right)!\left(\frac{2-k}{2}+m+\frac{1-l}{2}+n-1\right)!}{\left(\frac{2-k}{2}-m\right)!\left(\frac{1-l}{2}-n\right)!\left(\frac{2-k}{2}+m\right)!\left(\frac{1-l}{2}+n\right)!}R^{k+l,d}_{a+b,m+n}$
$+\frac{\kappa\,\Lambda}{2}\left((l-1)\,a-k\,b\right)\frac{\left(\frac{2-k}{2}-m+\frac{1-l}{2}-n\right)!\left(\frac{2-k}{2}+m+\frac{1-l}{2}+n\right)!}{\left(\frac{2-k}{2}-m\right)!\left(\frac{1-l}{2}-n\right)!\left(\frac{2-k}{2}+m\right)!\left(\frac{1-l}{2}+n\right)!}R^{k+l-2,d}_{a+b,m+n}\,.$ (40)
Upon the redefinition written in Eq.(11) and
$S^{q,d}_{a,m}=(q-m-1)!(q+m-1)!R^{3-2q,d}_{a,m}\,,$ (41)
the commutators (40) become
$\left[w_{a,m}^{p},S^{q,d}_{b,n}\right]=[m(q-1)-n(p-1)]S^{p+q-2,d}_{a+b,m+n}-\Lambda[a(q-1)-b(p-2)]S^{p+q-1,d}_{a+b,m+n}\,.$ (42)
While gravitational interactions are affected by curvature, we do not expect corrections to pure Yang-Mills thery, therefore the $S$-algebra of soft gauge currents should remain in its original form Strominger:2021lvk ,
$\left[S^{p,d}_{a,m},S^{q,e}_{b,n}\right]=-if^{deg}S^{p+q-1,g}_{a+b,m+n}\,.$ (43)
The soft currents $S^{1,d}_{0,0}$ satisfy
$\left[S^{1,d}_{0,0},S^{1,e}_{0,0}\right]=-if^{deg}S^{1,g}_{0,0}\,.$ (44)
and generate global gauge transformations. From Eq.(42), it follows that
$\left[L_{a},S^{1,d}_{0,0}\right]=\left[\bar{L}_{m},S^{1,d}_{0,0}\right]=\left[P_{k,l},S^{1,d}_{0,0}\right]=0\,.$ (45)
Although we do not have a solid argument supporting the cosmological deformation written in Eq.(42), it is easy to check that the full symmetry algebra of Einstein-Yang-Mills systems, written in Eqs.(15), (42) and (43), satisfies all Jacobi identities.
Discussion. — In this letter, we constructed a cosmological deformation of $\mathrm{w_{1+\infty}}$ algebra, by including commutator terms proportional to the cosmological constant. As a result, the Poincaré subalgebra was replaced by the symmetry algebra of dS or AdS, depending on the sign of the cosmological constant. It is striking that this deformation was extracted from the singularity structure of the graviton scattering amplitudes in spacetime with constant curvature, although a slight modification of the corresponding graviton OPEs was necessary to ensure a self-consistent algebra. It would be very interesting to uncover a deeper reason for this modification. A precise connection to the construction of Alday, Hansen and Silva Alday:2022uxp ; Alday:2022xwz ; Alday:2023jdk ; Alday:2023mvu remains to be understood. In particular, their higher-order curvature corrections contain higher order poles in Mandelstam variables. The physical interpretation of these poles is not clear.
It would be also interesting to see how the cosmological $w_{1+\infty}$ algebra is related to the asymptotic symmetries of dS spacetime, discussed in Refs.Strominger:2001pn ; Anninos:2010zf ; Anninos:2011jp ; Compere:2019bua ; Compere:2020lrt ; Compere:2023ktn ; Mao:2019ahc ; Bonga:2023eml . Furthermore, one might utilize the methods developed in Freidel:2021ytz to find a realization of the cosmological $w_{1+\infty}$ algebra on the gravitational phase space.
Acknowledgements
This work was inspired by a talk given by John Schwarz at the kickoff workshop of Simons Collaboration on Celestial Holography (Cambridge, MA, October 26-29, 2023), in which he proposed a modification of the OPEs similar to Eq.($\mathbf{w_{1+\infty}}$ Algebra with a Cosmological Constant                                    and the Celestial Sphere). We would like to thank John Schwarz and Andy Strominger for comments on the manuscript and to Fernando Alday for useful correspondence. TRT is supported by the National Science Foundation under Grants Number PHY-1913328 and PHY-2209903, by the NAWA Grant “Celestial Holography of Fundamental Interactions” and by the Simons Collaboration on Celestial Holography under grant MP-SCMPS-00001550-05. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation. BZ is supported by the Royal Society.

### 08 Dec 2023 v2

@@
 in the MHV helicity configuration,
-$A^{(1)}(s,t,u)=\kappa\Lambda\frac{\langle 12\rangle^{6}[43]}{\langle 13\rangle\langle 14\rangle\langle 23\rangle\langle 24\rangle\langle 34\rangle}\Big{(}\frac{1}{s}+\frac{1}{t}+\frac{1}{u}\Big{)},$ (1)
+$A^{(1)}(s,t,u)=\kappa^{2}\Lambda\frac{\langle 12\rangle^{6}[43]}{\langle 13\rangle\langle 14\rangle\langle 23\rangle\langle 24\rangle\langle 34\rangle}\Big{(}\frac{1}{s}+\frac{1}{t}+\frac{1}{u}\Big{)},$ (1)
 where $s,t,u$ are the Mandelstam variables and for the momentum spinors, we used the notation of Ref.Taylor:2017sph .
@@
 The leading collinear singularity has the form
-$A^{(1)}(s,t,u)\approx\kappa\Lambda\frac{\omega_{P}^{2}}{\omega_{3}^{2}\omega_{4}^{2}(z_{3}-z_{4})^{2}}\left(\frac{\langle 12\rangle^{6}}{\langle 1P\rangle^{2}\langle 2P\rangle^{2}}\right),$ (3)
+$A^{(1)}(s,t,u)\approx\kappa\Lambda\frac{\omega_{P}^{2}}{\omega_{3}^{2}\omega_{4}^{2}(z_{3}-z_{4})^{2}}\left(\kappa\frac{\langle 12\rangle^{6}}{\langle 1P\rangle^{2}\langle 2P\rangle^{2}}\right),$ (3)
 therefore we obtain the three-graviton amplitude (enclosed in the brackets) times the collinear factor with a double pole $(z_{3}-z_{4})^{-2}$.